Calculate the reaction free energy of

H2(g) + I2(g) 2HI(g)

when the concentrations are 0.006 mol L-1 (H2), 0.03 mol L-1 (I2), and 1.27 mol L-1 (HI), and the temperature is 700K. For this reaction Kc = 54 at 700 K.


+52,100 kJ mol-1

+52 kJ mol-1

-29.8 kJ mol-1

+29.8 kJ mol-1


1
Expert's answer
2020-09-21T06:50:25-0400

"\Delta""\mathring{G}" = RTlnKc

R = 8.314 j/kmol

T = 700 K

Kc = 54

"\frac{[HI][HI]}{[H2][I2]} =" "\frac{1.27][1.27]}{[0.006][0.03]}"

"=" 8960.55

"\Delta G =\Delta""\mathring{G}" + RTln"\theta"

"= -RTl"n(Kc) "+ RTl"n"\theta"

"= RTl"n"\frac{\theta}{Kc}"

"= 8.314 \times 700 \times l\frac{8960.55}{54}"

"= 29.8 kj\/mol"